A simple random sample of n = 21 people with early-onset dementia have the mean age at diagnosis of ¯x = 52.3 years and the standard deviation of s = 6.5 years.

(a) At 5% significance level perform a hypotheses test to determine whether mean age at diagnosis of all people with early-onset dementia is di↵erent from 55 years.

(b) Interpret your answer from part (a).
Answer #1

Solution :

\mu= 55

\bar x =52.3

S =6.5

n = 21

This is the two tailed test .

The null and alternative hypothesis is ,

H0 :  \mu  = 55

Ha : \mu    \neq 55

Test statistic = t

= (\bar x - \mu ) / S / \sqrt n

= (52.3 -55) / 6.5 / \sqrt 21

= −1.904

Test statistic = t = −1.904

P-value =0.0715

a ) \alpha = 0.05  

P-value ≥ \alpha

0.0715 ≥ 0.05

b ) Do not reject the null hypothesis .

There is insufficient evidence to suggest that